OpenStudy (anonymous):

3x+2 over (x-4)(x+4)(x-4)

OpenStudy (erinweeks):

are you looking for the x and y ?

OpenStudy (anonymous):

just x

OpenStudy (erinweeks):

ok

zepdrix (zepdrix):

What are you trying to do live? We have no equality sign, so we can't solve for x. Are you trying to simplify it? Or just expand out the brackets? :O

OpenStudy (anonymous):

yes trying just to simplify

OpenStudy (erinweeks):

this is what i got , i am confused on this question. -62 + 19x + 4x2 + -1x3 = 0

zepdrix (zepdrix):

\[\large \frac{3x+2}{(x-4)(x+4)(x-4)}\]Is this what the problem looks like?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Let's deal with just 2 of the brackets first, Notice how we have (x+4)(x-4) Those are called "conjugates", they're the SAME except with the opposite sign between them. Here is the rule for multiplying conjugates,\[\huge (a+b)(a-b)=a^2-b^2\]Applying this to our problem gives us,\[\huge (x+4)(x-4)=x^2-4^2\]Understand that part ok?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

That changes our equation to,\[\huge \frac{3x+2}{(x-4)(x^2-4^2)}\] Now to further expand, we just FOIL out the bottom.

OpenStudy (anonymous):

can we box it?, i never was taught the foil method

zepdrix (zepdrix):

box it? :) how does that work?

OpenStudy (anonymous):



zepdrix (zepdrix):

Oh cool :D

zepdrix (zepdrix):

So if we square the 4, we have, \[\huge (x-4)(x^2-16)\]and we can do the box method from here

OpenStudy (anonymous):

k

zepdrix (zepdrix):

Are you putting the square terms across the top? or along the side? I just wanna make sure I do the same :X

OpenStudy (anonymous):